The same amount of electric current was passed through two seperate electrolytic cells containing solutions of nickle nitrate and chromium nitrate respectively.If 0.3g of nickle was deposited in the first cell the amount of chromium deposited is?
 
 

Grade:12th pass

2 Answers

Arun
5 years ago
Dear student
 
mNi /mCr = eq mass of Ni/ eq mass of Cr
 
0.3/ mCr = 59/2 / 52/3
 
mCr = 0.176 gram
 
Regards
Arun (askIITians forum expert)
Mahima Kanawat
5 years ago
Dear student 
The no. Of equivalents of deposited ions will be same 
No. Of eq. = Wt * n factor / mol wt
Mol wt of Ni = 58
Mol wt of Cr =52
0.30*2/58 = wt of cr *2/52
Wt of deposited Cr = 0.26 gm
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT